I am stuck:
(x-2)^3=24
x^3-2^3=24
x^3-8=24
x^3=32
x=sqrt(32)
Is this correct?

Answer by stanbon(75887)

(x-2)^3=24
----
(x-2)^3 = 8*3
-----
x-2 = cbrt(8)*cbrt(3)
----
x = 2*cbrt(3) +2
=====================
Cheers,
Stan H.
=====================